OpenStudy (anonymous):

Click on all that are true about the graph of y > x^2 + 4x + 4. The parabola opens up. The line of symmetry is y = -2. The vertex is at (-2, 0). The parabola is dashed. The parabola is shaded outside.

OpenStudy (anonymous):

@satellite73 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

how far did you get?

OpenStudy (anonymous):

Umm... The first one and the last two..?

jimthompson5910 (jim_thompson5910):

are you able to graph y = x^2 + 4x + 4 ?

OpenStudy (anonymous):

I don't know... I don't have a graph

jimthompson5910 (jim_thompson5910):

do you have a graphing calculator?

OpenStudy (anonymous):

No...

jimthompson5910 (jim_thompson5910):

ok, here is what the graph looks like http://fooplot.com/plot/83f01cmra4

jimthompson5910 (jim_thompson5910):

you can use that website to graph other functions as well

jimthompson5910 (jim_thompson5910):

so where is the vertex on that blue parabolic curve?

OpenStudy (anonymous):

Okay I don't understand...

jimthompson5910 (jim_thompson5910):

are you able to see the graph/picture in that link I sent?

OpenStudy (anonymous):

Yes

jimthompson5910 (jim_thompson5910):

what does it look like?

OpenStudy (anonymous):

It opens up, it is on -2, it is solid, but I don't know where it is shaded at

jimthompson5910 (jim_thompson5910):

where is the lowest point as an ordered pair?

OpenStudy (anonymous):

What?

jimthompson5910 (jim_thompson5910):

for example, (-1,1) is on the graph but it is not the lowest point on this curve

jimthompson5910 (jim_thompson5910):

do you see how (-1,1) is a point on the parabola?

OpenStudy (anonymous):

No. I don't even know what a parabola is... I ask all my questions on here lol

OpenStudy (anonymous):

Is it the first three answers?

jimthompson5910 (jim_thompson5910):

a parabola is that bowl-shaped curve

OpenStudy (anonymous):

Alright

jimthompson5910 (jim_thompson5910):

the lowest point (when the parabola opens upward) is the vertex

jimthompson5910 (jim_thompson5910):

what is the vertex of this parabola?

OpenStudy (anonymous):

-2 ?

jimthompson5910 (jim_thompson5910):

it has to be an ordered pair

jimthompson5910 (jim_thompson5910):

in the form (x,y)

OpenStudy (anonymous):

(-2, 0) ?

jimthompson5910 (jim_thompson5910):

much better

jimthompson5910 (jim_thompson5910):

you start at (0,0) and you go 2 units to the left since x =-2 and you don't move along the y axis since y = 0

jimthompson5910 (jim_thompson5910):

The line of symmetry is y = -2. that is false because the axis of symmetry is really x = -2

jimthompson5910 (jim_thompson5910):

The statement "The line of symmetry is y = -2." would only be true if the parabola opened up sideways, but it does not.

jimthompson5910 (jim_thompson5910):

So that covers the equation y = x^2 + 4x + 4 to graph y > x^2 + 4x + 4, we simply graph y = x^2 + 4x + 4, then we follow these steps step 1) make the parabola a dashed line (because we do NOT include the boundary) step 2) shade above the parabola (ie on the inside of it)

jimthompson5910 (jim_thompson5910):

so "The parabola is dashed." is true but "The parabola is shaded outside." is false

OpenStudy (anonymous):

So it is the first, third, and fourth answers?

jimthompson5910 (jim_thompson5910):

That is correct.

OpenStudy (anonymous):

Thank you:) I have some other questions to post. Would you mind looking at them?

jimthompson5910 (jim_thompson5910):

I'll help with 2 more